# Hierarchical Tri-manual Planning for Vision-assisted Fruit Harvesting with Quadrupedal Robots

### v1

@@ VI-B Outdoor Testing
We evaluate the performance of visual detection and harvesting of real avocados in natural environments. After teleoperating the robot to approach the avocado tree, the robot initiates the NBV planning and follows the generated trajectory to ensure that the target avocados fall within the feasible subset of the robot’s reachable workspace. Upon reaching the desired pose from the NBV planner, the learning-based visual perception system is activated using the RGB-D camera to perform target detection and pose estimation. If the target is detected outside the dual-arm system’s workspace, a message will be sent to the operator, prompting them to teleoperate the robot and re-initiate the framework.
Fig. 8 illustrates a sequence of successfully retrieving an avocado. We also run 10 trials and recorded the success rate in a single attempt. Table I contains the results. We observe a drop in the success rate in the natural environment compared to the controlled laboratory settings. Several factors may account for this difference. The pose estimation provided by the camera is subject to noise, which is further exacerbated by the movements of the Spot robot and its arm. Ambient wind in the orchard causes the target avocados to shake, presenting additional challenges for both visual detection and precise grasping. However, the success rate test only accounts for a single harvesting attempt, and the overall efficiency can be significantly improved if multiple attempts are considered and re-run the framework.

### v2

@@ VI-B Outdoor Testing
We evaluate the performance of visual detection and harvesting of real avocados in natural environments. After teleoperating the robot to approach the avocado tree, the robot initiates the NBV planning and follows the generated trajectory to ensure that the target avocados fall within the feasible subset of the robot’s reachable workspace. Upon reaching the desired pose from the NBV planner, the learning-based visual perception system is activated using the RGB-D camera to perform target detection and pose estimation. If the target is detected outside the dual-arm system’s workspace, a message is sent to the operator, prompting to teleoperate the robot and re-initiate the framework.
Figure 8 illustrates a sequence of successfully retrieving an avocado. We also run 10 trials and recorded the success rate in a single attempt. Table I contains the results. We observe a drop in the success rate in the natural environment compared to the controlled laboratory settings. Several factors may account for this difference. The pose estimation provided by the camera is subject to noise, which is further exacerbated by the movements of the Spot robot and its arm. Ambient wind in the orchard causes the target avocados to shake, presenting additional challenges for both visual detection and precise grasping. However, the success rate test only accounts for a single harvesting attempt, and the overall efficiency can be significantly improved if multiple attempts are considered and re-run the framework.